OpenStudy (anonymous):

Find the value of the following: -3^2 + ( (5 (8-6) ) / 2 + 45 ) / 5

OpenStudy (cwrw238):

this is an exercise in order of operations have you heard of PEDMAS?

OpenStudy (cwrw238):

its an acronym for order of operations Parentheses Exponents Divide Multiply Add Subtract

OpenStudy (anonymous):

Ye, I am familiar with it. It has been a long time since I have solved an equation like this. The answer I got is 19. I wanted to check my work.

OpenStudy (cwrw238):

well first work out the parentheses from the inside out so we have 8-6 = 2 then 5 * 2 = 10 10 / 2 = 5 5 + 45 = 50 50 / 5 = 10

OpenStudy (cwrw238):

the -3^2 is ambiguous it might mean 9 or -9 i think its -9 because if the -3 is to be squared it should be (-3)^2

OpenStudy (cwrw238):

so this gives -9+ 10 = 1

OpenStudy (cwrw238):

-3^2 is a subtraction and an exponent - the exponent is processed first

OpenStudy (anonymous):

Thank you :)